### v1

@@ 3. Regularity of self-similar measures
where $C_{k}$ can be formulated in terms of $\sigma,\varepsilon,c$. We argue by induction on $k$. The case $k=0$ is clear because $\sigma$ has a moment of positive order (Lemma˜3.1).
We assume now the result holds for $1,\dots,k-1$, and establish it for $k$. We write $\tau_{\varepsilon}:(\operatorname{Sim}(\mathbb{R}^{d})^{+})^{\mathbb{N}}\rightarrow\mathbb{N}\cup\{\infty\}$ the stopping time defined by
$\tau_{\varepsilon}({\underline{\phi}})=\inf\{n\geq 1\,:\,\mathtt{r}_{\phi_{1}\dots\phi_{n}}<\varepsilon\}$
As $\lambda$ is contracting in average, we have that $\tau_{\varepsilon}$ is $\lambda^{\mathbb{N}}$-almost surely finite. We set $\lambda^{*\tau_{\varepsilon}}$ the distribution of $\phi_{1}\dots\phi_{\tau_{\varepsilon}({\underline{\phi}})}$ as ${\underline{\phi}}\sim\lambda^{\mathbb{N}}$. It is known that $\sigma$ is $\lambda^{*\tau_{\varepsilon}}$-stationary, see e.g [3, Lemma A.2]. By the finite exponential moment of $\lambda$, it can be shown that the variable $\varepsilon/\mathtt{r}_{\phi}$ where $\phi\sim\lambda^{*\tau_{\varepsilon}}$ has a moment of positive order independently of $\varepsilon$: there exists $\gamma=\gamma(\sigma)>0$ such that $\sup_{\varepsilon>0}\int(\varepsilon/\mathtt{r}_{\phi})^{\gamma}\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)<2$, see e.g. the proof of [4, Proposition A.18].
Recalling the notation $\mathtt{b}_{\phi}=\phi(0)$. Since $\lambda$ is strongly irreducible, $\sigma$ gives zero mass to proper affine subspaces. Combined with Lemmas 2.1, 3.1, this allows to introduce (small enough) $\delta,\gamma^{\prime}\in(0,1/2)$ depending on $\sigma$ only, and such that
$\sup_{\varepsilon\leq\delta}\sup_{{\mathcal{L}}}\lambda^{*\tau_{\varepsilon}}\{\phi\,:\,\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})\leq\delta\}<\frac{1}{10^{4}},$
and for all $T\geq\delta^{-1}/2$,
$\sup_{\varepsilon\leq\delta}\lambda^{*\tau_{\varepsilon}}\{\phi\,:\,\|\mathtt{b}_{\phi}\|\geq T\}<\frac{1}{10^{4}}(1+T)^{-\gamma^{\prime}}.$
Let $s\in[\varepsilon^{k+1},\varepsilon^{k})$, $T\geq 0$, and consider $\mathcal{L}$} such that $\operatorname{dist}(0,{\mathcal{L}})\geq T$. Set $T^{\prime}=\delta^{2}\max(1,T)$. Using $\lambda^{*\tau_{\varepsilon}}*\sigma=\sigma$ and distinguishing according to the position of $\mathtt{b}_{\phi}$, we have
$\sigma({\mathcal{L}}^{(s)})=\underbrace{\int_{\{\phi\,:\,\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})\geq T^{\prime}\}}\phi_{\star}\sigma({\mathcal{L}}^{(s)})\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)}_{A}\,\,\,+\,\,\,\underbrace{\int_{\{\phi\,:\,\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})<T^{\prime}\}}\phi_{\star}\sigma({\mathcal{L}}^{(s)})\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)}_{B}.$
Using the induction hypothesis, we establish an upper bound for $A$. More precisely, noting that $\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})\geq T^{\prime}$ implies $\operatorname{dist}(0,{\phi}^{-1}{\mathcal{L}})\geq\mathtt{r}_{\phi}^{-1}T^{\prime}$, then using that $\mathtt{r}_{\phi}^{-1}s\geq\varepsilon^{k}$ for all $\phi\in\operatorname{supp}\lambda^{*\tau_{\varepsilon}}$ to apply the induction hypothesis, we find
$A$ $\leq\int_{\operatorname{Sim}(\mathbb{R}^{d})^{+}}I_{\mathtt{r}_{\phi}^{-1}s,\mathtt{r}_{\phi}^{-1}T^{\prime}}\,\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)$
$\leq C_{k-1}\int_{\operatorname{Sim}(\mathbb{R}^{d})^{+}}(\mathtt{r}_{\phi}^{-1}s(1+\mathtt{r}_{\phi}^{-1}T^{\prime})^{-1})^{c}\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)$
$\leq C_{k-1}2\delta^{-2c}s^{c}(1+T)^{-c}.$
We now bound $B$. We have
@@
$B$ $\leq\int_{\{\phi\,:\,\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})<T^{\prime}\}}I_{\mathtt{r}_{\phi}^{-1}s,0}\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)$
@@
$\leq C_{k-1}s^{c}\varepsilon^{-c}\int_{\{\phi\,:\,\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})<T^{\prime}\}}(\varepsilon/\mathtt{r}_{\phi})^{c}\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)$
$\leq C_{k-1}s^{c}\varepsilon^{-c}2\sqrt{\lambda^{*\tau_{\varepsilon}}\{\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})<T^{\prime}\}}$
where the last inequality relies on Cauchy-Schwarz, and assumes $c\leq\gamma/2$ to guarantee $\int_{\operatorname{Sim}(\mathbb{R}^{d})^{+}}(\varepsilon/\mathtt{r}_{\phi})^{2c}\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)\leq 2$. Note also, by distinguishing the cases $T^{\prime}\leq\delta$ (i.e. $T\leq\delta^{-1}$) and $T^{\prime}>\delta$ (i.e. $T>\delta^{-1}$), that the definition of $\delta$ yields
@@
$\sqrt{\lambda^{*\tau_{\varepsilon}}\{\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})<T^{\prime}\}}\leq\frac{1}{50}(1+\delta T)^{-\gamma^{\prime}/2}.$
Finally
$B\leq C_{k-1}s^{c}\varepsilon^{-c}\frac{1}{25}(1+\delta T)^{-\gamma^{\prime}/2}.$
From the above, we have obtained
$I_{s,T}\leq C^{\prime}s^{c}(1+T)^{-c}$
where
$C^{\prime}$ $=2(\delta^{-2c}+\varepsilon^{-c}\frac{1}{50}(1+T)^{c}(1+\delta T)^{-\gamma^{\prime}/2})C_{k-1}$
$\leq 2\delta^{-2c}(1+\varepsilon^{-c}\frac{1}{50})C_{k-1}$
where the second inequality assumed $c<\gamma^{\prime}/2$. We now check that $c,\delta$ and $\varepsilon$ can in fact be chosen so that
$2\delta^{-2c}(1+\varepsilon^{-c}\frac{1}{50})\leq\varepsilon^{-c/2}.$
Indeed, it suffices to select $c\lll_{\delta}1$ such that $2\delta^{-2c}<21/10$, then $\varepsilon$ such that $\varepsilon^{c/2}=2/5$. We have thus established
$C^{\prime}\leq\varepsilon^{-c/2}C_{k-1}.$
The above justifies (3.1), with constants $C_{k}$ given by $C_{k}=\varepsilon^{-kc/2}C_{0}$ where $C_{0}=C_{0}(\sigma)>1$. It follows that for every $s>0,T\geq 0$,
$I_{s,T}\leq C^{\prime}_{0}s^{c/3}(1+T)^{-c},$
where $C_{0}^{\prime}=C_{0}\varepsilon^{-c}$ is bounded depending on $\sigma$ only, and the proof of the proposition is complete. ∎
@@
We upgrade Proposition˜3.2 into a Hölder control on the concentration of $\sigma$ near algebraic subvarieties of $\mathbb{R}^{d}$. Given $l\in\mathbb{N}$, we set $\mathcal{P}_{d,l}$ the vector space of real polynomial functions on $\mathbb{R}^{d}$ of degree at most $l$. We equip $\mathcal{P}_{d,l}$ with the supremum norm on the coefficients, which we write $\|\cdot\|$.
Theorem 3.3 (Non-concentration near subvarieties).
For every $l\in\mathbb{N}$, $P\in\mathcal{P}_{d,l}$ with $\|P\|\geq 1$, and $\varepsilon>0$, we have
@@
$\sigma(\boldsymbol{s}\in\mathbb{R}^{d}\,:\,|P(\boldsymbol{s})|\leq\varepsilon)\leq C\varepsilon^{c},$
@@
The idea of the proof of ˜3.3 is to use the self-similarity of $\sigma$ to write $\sigma$ as a convex combination of measures $(\sigma_{j})_{j}$ obtained from $\sigma$ by pushing via affine maps, and with each $\sigma_{j}$ living at scale $\varepsilon^{3/4}$, say roughly supported in $B(x_{j},\varepsilon^{3/4})$. For each $\sigma_{j}$, we may approximate the polynomial map $P$ by its Taylor expansion up to order $1$ at $x_{j}$, and apply non-concentration near affine hyperplanes to deduce the estimate in ˜3.3. In fact, to make this argument work, we also need to make sure that most of the $x_{j}$’s are located where the gradient $\nabla P$ of $P$ is not too small. But $\nabla P$ is polynomial of smaller degree, and the distribution of the $x_{j}$’s resembles that of $\sigma$, whence we may guarantee this using an inductive approach. A related strategy is exploited in [33, Section 7] in the context of absolutely decaying measures.
@@
In the same manner, one can show that $\sigma$ is not concentrated near submanifolds $M$ of $\mathbb{R}^{d}$ such that $\dim M<d$, as long as $M$ is not too badly approximated by its tangent subspaces (e.g. if $\exp_{x}:B^{T_{x}M}_{1}\rightarrow M$ has uniformly bounded order $2$ derivatives for all $x$ in the manifold).
Proof of ˜3.3.
@@
$E_{\varepsilon}:=\{\boldsymbol{s}\in\mathbb{R}^{d}:|P(\boldsymbol{s})|\leq\varepsilon\}.$
@@
$F_{\delta}:=\{\boldsymbol{s}\in\mathbb{R}^{d}:\|\nabla P(\boldsymbol{s})\|\leq\delta\}$
@@
We decompose $\sigma$ as a combination of measures living at scale $\varepsilon^{3/4}$, and group them into 3 categories, distinguishing whether they are centered around a point outside of $B_{R}$, or within $B_{R}$, and in the second case whether the point is in $F_{\delta}$ or not. More formally, recalling the Lyapunov exponent $\ell$ from (2.1), we set $n=\lfloor\frac{3}{4\ell}\lvert\log\varepsilon\rvert\rfloor$, so that for $\phi\sim\lambda^{*n}$, we have $r_{\phi}$ close to $\varepsilon^{3/4}$. We then have by $\lambda$-stationarity of $\sigma$:
$\sigma(E_{\varepsilon})$ $=\int_{\operatorname{Sim}(\mathbb{R}^{d})}\phi_{\star}\sigma(E_{\varepsilon})\,\mathrm{d}\lambda^{*n}(\phi)=I_{1}+I_{2}+I_{3}$
@@
$I_{1}=\int_{\phi(0)\notin B_{R}}\phi_{\star}\sigma(E_{\varepsilon})\,\mathrm{d}\lambda^{*n}(\phi),\,\,\,\,\,\,\,\,I_{2}=\int_{\phi(0)\in B_{R}\cap F_{\delta}}\phi_{\star}\sigma(E_{\varepsilon})\,\mathrm{d}\lambda^{*n}(\phi),$
$I_{3}=\int_{\phi(0)\in B_{R}\smallsetminus F_{\delta}}\phi_{\star}\sigma(E_{\varepsilon})\,\mathrm{d}\lambda^{*n}(\phi).$
Combining Lemma˜3.1 and Lemma˜2.1, we have for some $\gamma=\gamma(\lambda)>0$,
(3.2) $I_{1}\ll R^{-\gamma}.$
@@
(3.3) $\|\nabla P\|\gg_{l}R^{-(1+l)}.$
To see why, note first that $\|P\|=1$ implies $\sup_{B_{1}}|P|\geq\eta$ where $\eta=\eta(d,l)>0$. Now if $\|\nabla P\|_{B_{R}}\leq R^{-1}\eta/4$, we get $\inf_{B_{R}}|P|\geq\eta/2$. In this scenario, provided $\varepsilon\leq\eta/4$, we have $E_{\varepsilon}\subseteq\mathbb{R}^{d}\smallsetminus B_{R}$, and the result follows from the finite moment of $\sigma$ (Lemma˜3.1). Hence we may suppose $\|\nabla P\|_{B_{R}}>R^{-1}\eta/4$, and (3.3) follows by passing to the norm $\|\cdot\|$.
Let $c_{1}>0$. Note that the upper bound $\|\nabla P\|_{B_{R+1}}\ll_{l}R^{l}$ implies that the $\varepsilon^{c_{1}}$-neighborhood of $B_{R}\cap F_{\delta}$ is included in $F_{\delta+O_{l}(\varepsilon^{c_{1}}R^{l})}$. Choosing $c_{1}=c_{1}(\sigma)$ small enough, we deduce from Lemma˜2.1 that
$I_{2}$ $\leq\sigma^{(n)}(B_{R}\cap F_{\delta})\leq\sigma(F_{\delta+O_{l}(\varepsilon^{c_{1}}R^{l})})+O(\varepsilon^{c_{1}}).$
@@
(3.4) $I_{2}\ll_{l}R^{(1+l)c_{2}}(\delta+\varepsilon^{c_{1}}R^{l})^{c_{2}}+\varepsilon^{c_{1}}$

### v2

@@ 4.1. Hölder regularity near affine subspaces
We argue by induction on $k$. The case $k=0$ is clear because $\sigma$ has a moment of positive order (Lemma˜3.1) and we are allowed to pick $c>0$ small compared to the constant $\gamma$ in Lemma˜3.1.
Assume (4.4) is true for some $k\in\mathbb{N}$ and let $s\geq\varepsilon^{k+1}$. Note that by the $\lambda^{*\tau_{\varepsilon}}$-stationarity of $\sigma$ (Lemma˜4.3 (1)), we have
$\sigma({\mathcal{L}}^{(s)})=\int\sigma\bigl(\phi^{-1}({\mathcal{L}}^{(s)})\bigr)\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi).$
Since $\phi$ is a similarity of contraction ratio $\mathtt{r}_{\phi}$, we have $\phi^{-1}({\mathcal{L}}^{(s)})=(\phi^{-1}{\mathcal{L}})^{(\mathtt{r}_{\phi}^{-1}s)}$ as well as $\mathtt{r}_{\phi}\operatorname{dist}(0,\phi^{-1}{\mathcal{L}})=\operatorname{dist}(\phi(0),{\mathcal{L}})=\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})$. By the definition of the stopping time, for $\lambda^{*\tau_{\varepsilon}}$-almost every $\phi$, we have $\mathtt{r}_{\phi}^{-1}s\geq\varepsilon^{k}$, therefore we may invoke the induction hypothesis to deduce
$\sigma({\mathcal{L}}^{(s)})$ $=\int\sigma((\phi^{-1}{\mathcal{L}})^{(\mathtt{r}_{\phi}^{-1}s)})\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)$
@@
$\leq C_{k}\int(\mathtt{r}_{\phi}^{-1}s)^{c}\bigl(1+\operatorname{dist}(0,\phi^{-1}{\mathcal{L}})\bigr)^{-c}\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)$
@@
$=C_{k}s^{c}\int\bigl(\mathtt{r}_{\phi}+\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})\bigr)^{-c}\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi).$ (4.5)
Note that in (4.5), the integral is bounded by $O(\varepsilon^{c})$ (by ignoring the term $\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})$ and applying (4.1)), but this is not good enough to conclude, as it would yield $C_{k+1}\geq\varepsilon^{-c}C_{k}$, so the sequence $C_{k}$ would grow too fast. To upper bound the integral in (4.5), we need to exploit the contribution of $\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})$.
Let $\delta=\delta(\lambda)\in(0,1/10)$ be as in Lemma˜4.3, write $T:=\operatorname{dist}(0,{\mathcal{L}})$ for a shorthand and set
$A:=\bigl\{\,\phi\in\operatorname{Sim}(\mathbb{R}^{d})\,:\,\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})\leq\delta^{2}(1+T)\,\bigr\}.$
Using
$\mathtt{r}_{\phi}+\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})\geq\begin{cases}\mathtt{r}_{\phi}&\text{ if }\phi\in A,\\ \delta^{2}(1+T)&\text{ otherwise,}\end{cases}$
and (4.5), followed by the Cauchy-Schwarz inequality and (4.1), we obtain
@@
$\frac{\sigma({\mathcal{L}}^{(s)})}{C_{k}s^{c}}$ $\leq\int_{A}\mathtt{r}_{\phi}^{-c}\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)+\delta^{-2c}(1+T)^{-c}$
$\leq 2\varepsilon^{-c}\lambda^{*\tau_{\varepsilon}}(A)^{1/2}+\delta^{-2c}(1+T)^{-c}.$ (4.6)
We claim that
$\lambda^{*\tau_{\varepsilon}}(A)\leq 2^{-8}(1+T)^{-2c},$ (4.7)
provided $c$ is small enough depending on $\lambda$. Assuming this claim, one can conclude the proof as follows: Equations (4.6), (4.7) together imply
$\sigma({\mathcal{L}}^{(s)})\leq\bigl(8^{-1}\varepsilon^{-c}+\delta^{-2c}\bigr)C_{k}s^{c}(1+T)^{-c}.$
Choosing $c$ small so that $\delta^{-2c}\leq 5/4$, and $\varepsilon=\varepsilon(c)\in(0,1)$ so that $\varepsilon^{-c}=4$, we find $8^{-1}\varepsilon^{-c}+\delta^{-2c}\leq\varepsilon^{-c/2}$, and this finishes the proof of the induction step for (4.4).
It remains to show the claim (4.7). To do so, note that if $\delta^{2}(1+T)\leq\delta$, that is if $T\leq\delta^{-1}-1$, then by (4.2), $\lambda^{*\tau_{\varepsilon}}(A)\leq 2^{-9}$. For such $T$, we can ensure $2^{-1}\leq(1+T)^{-2c}$ by asking $c\lll 1$ (depending on $\delta=\delta(\lambda)$). In the other case where $T>\delta^{-1}-1\geq\delta^{-1}/2$, note that the triangle inequality implies for all $\phi\in A$,
$\lVert\mathtt{b}_{\phi}\rVert\geq\operatorname{dist}(0,{\mathcal{L}})-\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})\geq T-\delta^{2}(1+T)\geq T/2.$
In this case, we conclude using (4.3): for $c\lll 1$,
$\lambda^{*\tau_{\varepsilon}}(A)\leq\lambda^{*\tau_{\varepsilon}}\{\,\phi\,:\,\lVert\mathtt{b}_{\phi}\rVert\geq T/2\,\}\leq 2^{-8}(1+T)^{-2c}.\qed$
4.2. Hölder regularity near subvarieties
@@
We upgrade Proposition˜4.2 into a Hölder control on the concentration of $\sigma$ near algebraic subvarieties of $\mathbb{R}^{d}$. Given $l\in\mathbb{N}$, we set $\mathcal{P}_{d,l}$ the vector space of real polynomial functions on $\mathbb{R}^{d}$ of degree at most $l$. We equip $\mathcal{P}_{d,l}$ with the supremum norm on the coefficients, which we write $\|\cdot\|$.
Theorem 4.4 (Hölder regularity near subvarieties).
For every $l\in\mathbb{N}$, $P\in\mathcal{P}_{d,l}$ and $\varepsilon>0$, we have
@@
$\sigma\bigl\{\,\boldsymbol{s}\in\mathbb{R}^{d}\,:\,\lvert P(\boldsymbol{s})\rvert\leq\varepsilon\lVert P\rVert\,\bigr\}\leq C\varepsilon^{c},$
@@
The idea of the proof of ˜4.4 is to use the self-similarity of $\sigma$ to write $\sigma$ as a convex combination of measures $(\sigma_{j})_{j}$ obtained from $\sigma$ by pushing via similarities, and with each $\sigma_{j}$ living at scale $\varepsilon^{3/4}$, say roughly supported in $B^{\mathbb{R}^{d}}_{\varepsilon^{3/4}}+x_{j}$ for some $x_{j}$. For each $\sigma_{j}$, we may approximate the polynomial map $P$ by its Taylor expansion up to order $1$ at $x_{j}$, and apply non-concentration near affine hyperplanes to deduce the estimate in ˜4.4. In fact, to make this argument work, we also need to make sure that most of the $x_{j}$’s are located where the gradient $\nabla P$ of $P$ is not too small. But $\nabla P$ is a polynomial of smaller degree, and the distribution of the $x_{j}$’s resembles that of $\sigma$, whence we may guarantee this using an inductive approach. A related strategy is exploited in [43, Section 7] in the context of absolutely decaying measures.
@@
In the same manner, one can show that $\sigma$ is not concentrated near submanifolds $M$ of $\mathbb{R}^{d}$ such that $\dim M<d$, as long as $M$ is not too badly approximated by its tangent subspaces (e.g. if $\exp_{x}:B^{T_{x}M}_{1}\rightarrow M$ has uniformly bounded order $2$ derivatives for all $x$ in the manifold $M$).
Proof of ˜4.4.
@@
$E_{\varepsilon}:=\bigl\{\,\boldsymbol{s}\in\mathbb{R}^{d}:|P(\boldsymbol{s})|\leq\varepsilon\,\bigr\}.$
@@
$F_{\delta}:=\bigl\{\,\boldsymbol{s}\in\mathbb{R}^{d}:\|\nabla P(\boldsymbol{s})\|\leq\delta\,\bigr\}$
@@
We decompose $\sigma$ as a combination of measures living at scale $\varepsilon^{3/4}$, and group them into 3 categories, distinguishing whether they are centered around a point outside of $B_{R}$, or within $B_{R}$, and in the second case whether the point is in $F_{\delta}$ or not. More precisely, recalling the Lyapunov exponent $\ell$ from (2.1), we set $n=\lfloor\frac{3}{4\ell}\lvert\log\varepsilon\rvert\rfloor$, so that for $\phi\sim\lambda^{*n}$, we have $\mathtt{r}_{\phi}$ close to $\varepsilon^{3/4}$ most of the time. We then have by the $\lambda$-stationarity of $\sigma$:
$\sigma(E_{\varepsilon})=\int_{\operatorname{Sim}(\mathbb{R}^{d})}\phi_{\star}\sigma(E_{\varepsilon})\,\mathrm{d}\lambda^{*n}(\phi).$
Decomposing the integral depending on the location of $\mathtt{b}_{\phi}$, we obtain
$\sigma(E_{\varepsilon})\leq I_{1}+I_{2}+I_{3},$
@@
$I_{1}=\lambda^{*n}\{\phi\,:\,\mathtt{b}_{\phi}\notin B_{R}\},\quad I_{2}=\lambda^{*n}\{\phi\,:\,\mathtt{b}_{\phi}\in B_{R}\cap F_{\delta}\}\quad\text{and}$
$I_{3}=\int_{\mathtt{b}_{\phi}\notin F_{\delta}}\phi_{\star}\sigma(E_{\varepsilon})\,\mathrm{d}\lambda^{*n}(\phi).$
By Lemma˜3.3, we have for some $\gamma=\gamma(\lambda)>0$,
$I_{1}\ll R^{-\gamma}.$ (4.8)
@@
$\|\nabla P\|\gg_{l}R^{-(1+l)}.$ (4.9)
To see why, note first that $\|P\|=1$ implies $\sup_{B_{1}}|P|\geq\eta$ where $\eta=\eta(d,l)>0$. Now if $\|\nabla P\|_{B_{R}}\leq R^{-1}\eta/4$, we would have $\inf_{B_{R}}|P|\geq\eta/2$. Provided moreover that $\varepsilon\leq\eta/4$, we would have $E_{\varepsilon}\subseteq\mathbb{R}^{d}\smallsetminus B_{R}$, and the desired upper bound on $\sigma(E_{\varepsilon})$ follows from the finite moment of $\sigma$ (Lemma˜3.1). Hence we may suppose without loss of generality that $\lVert\nabla P\rVert_{B_{R}}>R^{-1}\eta/4$, and (4.9) follows by passing to the norm $\lVert\cdot\rVert$.
Let $c_{1}>0$. Note that the upper bound $\lVert\nabla P\rVert_{B_{R+1}}\ll_{l}R^{l}$ implies that the $\varepsilon^{c_{1}}$-neighborhood of $B_{R}\cap F_{\delta}$ is included in $F_{\delta+O_{l}(\varepsilon^{c_{1}}R^{l})}$. Choosing $c_{1}=c_{1}(\sigma)$ small enough, we deduce from Lemma˜3.2 that
$I_{2}$ $\leq\lambda^{*n}*\delta_{0}(B_{R}\cap F_{\delta})\leq\sigma(F_{\delta+O_{l}(\varepsilon^{c_{1}}R^{l})})+O(\varepsilon^{c_{1}}).$
@@
$I_{2}\ll_{l}R^{(1+l)c_{2}}(\delta+\varepsilon^{c_{1}}R^{l})^{c_{2}}+\varepsilon^{c_{1}}$ (4.10)